OpenStudy (melodious):

HELP!!! ^_^ Solve x and y 1/2(3x+2y) + 5/3(3x-2y)= -3/2 5/4(3x+2y) - 3/5(3x-2y)= 61/60

OpenStudy (robtobey2):

\[{x = -(1/9), y = 1/3}\]

OpenStudy (melodious):

How do u do it? Plzz

OpenStudy (robtobey2):

Expand out the products and then get rid of the fractions. \[6 \left(\frac{13 x}{2}-\frac{7 y}{3}\right)=\frac{1}{2} (-3) 6\to 39 x-14 y=-9 \]Do the same with the second equation. Then solve for x and y.

OpenStudy (usukidoll):

@robtobey2 I think you're needed back here.

OpenStudy (aaronandyson):

Do you still need help?

OpenStudy (mathmate):

@robtobey2 http://prntscr.com/bbibl0

OpenStudy (melodious):

HELP is NEEDED !!!

OpenStudy (mathmate):

@melodious "Help is needed! " is not specific. Please post where do you need help on, or the part that you do not understand, so people can help you even when you're not online.

OpenStudy (melodious):

I just want the steps to reach to the final answer @mathmate

OpenStudy (mathmate):

Let me take a look! brb.

OpenStudy (mathmate):

@melodious Here's how it goes. Step 1: Multiply out each of the equations and simplify it to the form Ax+by=C For example, the first equation is: (1/2)(3x+2y)+(5/3)(3x-2y)=-3/2 Multiply both sides by 6 (common denominator of all the fractions, don't forget to multiply every term, and the negative signs: 6(1/2)(3x+2y)+6(5/3)(3x-2y)=6(-3/2) 3(3x+2y)+10(3x-2y)=-9 9x+6y+30x-20y=-9 39x-14y=-9 I will let you simplify the second equation. After that, it will become a system of 2 liinear equations which can be solved readily by elimination. So please proceed to simplify and solve, and if you get stuck, please post what you've got, and where you get stuck. Good luck!